B) Hoaa expands its factory in Marysville, Ohio. 7. An economy produces the following goods: cars, textbooks, hamburgers, shirts and cotton. The quantities (Q) and prices (P) for the years of 2005 and 2010 are as follows 2 points each Q in 2005 P in 2005 Q in 2010 P in 2010 Textbooks 90 Hamburgers 75 Shirts Cotton 50 $2 $30 100 100 50 8,000 $60 $2 $25 $.6 50 10,000$.8 (a) Calculate nominal GDP for the 2 years. (b) What is the percent change in nominal GDP from year 2005 to 2010. (c) Calculate real GDP using 2005 as the base year for both years. (d) What is the percent change in real GDP from year 2005 to 2010 using 2005 as the base year. (e) Why are the growth rates calculated in 7b and 7d different? Which one more accurately reflects changes in output? (t) Compute the GDP deflator for both years. (g) Using the GDP deflator, find the inflation rate. (h) List one difference between the GDP deflator and the Consumer Price Index as measures of general price levels in the economy
Answer #1

7.a) Nominal GDP is calculated by multiplying current year price with current year quantity.

Nominal GDP in 2005 = (price of textbooks in 2005 * quantity of textbooks in 2005) + (price of hamburgers in 2005 * quantity of hamburgers in 2005) + ( price of shirts in 2005 * quantitity of shirts in 2005) + (price of cotton in 2005 * quantity of cotton in 2005) = (50*90)+(2*75)+(30*50)+(0.8*10,000)=$(4500 + 150+ 1500 + 8000) = $14,150

NGDP in 2010 = (60*100) + (2*100)+(25*50)+(0.6*8000) = $(6000 + 200+ 1250 + 4800) = $12,250

B) % change in NGDP from year 2005 to 2010 = [(12,250 - 14,150)/14,150]*100 = -13.43%

It means from 2005 to 2010, NGDP decreases by 13.43%

C) Real GDP = quantity of good in current year * price of that good in base year.

In 2010, Real GDP = (price of textbooks in 2005 * quantity in 2010) + (price of hamburgers in 2005 * quantity in 2010) + (price of shirts in 2005 * quantity in 2010) + (price of cotton in 2005*quantity in 2010) = (50*100) + (2*100) + (30*50) + (0.8 * 8000) = $13,100

In 2005, RGDP = (50*90)+(2*75)+(30*50)+(0.8*10000) = $14,150

D) % change in RGDP from the year 2005 to 2010 = [(13,100 - 14,150)/14,150]*100 = - 7.42%

That means RGDP decreases by 7.42% from the year 2005 to 2010.